OpenStudy (loser66):

find an integrating factor and solve the given equation \[1+(\frac{x}{y}-siny)y'=0\] Please, help

OpenStudy (anonymous):

\[1+\left( \frac{ x }{ y }-\sin y \right)y \prime=0\] \[\left( \frac{ x }{ y }-\sin y \right)\frac{ dy }{ dx }=-1\] \[\frac{ x }{ y }-\sin y=-\frac{ dx }{dy }\] \[\frac{ dx }{ dy }+\frac{ 1 }{ y }x=\sin y\] \[I.F.=e ^{\int\limits \frac{ 1 }{ y }dy}=e ^{\ln y}=y\] (assuming y>0) \[C.S. is x y=\int\limits y \sin y dy +c\] \[xy=y (-\cos y)-\int\limits 1*\left(- \cos y \right)dy+c\] \[xy=-y \cos y+\sin y+c\]

OpenStudy (loser66):

Thanks for reply and sorry for not being here to get help. The net was so bad here. I saw you wrote something but couldn't access to the net. Again, thank you very much. I appreciate.

OpenStudy (anonymous):

yw